ANALYSIS OF THE SPECTRUM OF ETHYLENE IN THE 800–1500 CM^{−1} REGION USING TENSORIAL FORMALISM: FREQUENCIES AND INTENSITIES

\underline{M. ROTGER}, V. BOUDON, Laboratoire de Physique de l’Université de Bourgogne, UMR CNRS 5027, 9, av. Alain Savary, B.P. 47870, F-21078 Dijon Cedex, France; A. FAYT, Molecular spectroscopy laboratory, Université Catholique de Louvain, Chemin du Cyclotron 2, B-1348 Louvain-La-Neuve, Belgium; J. VANDER AUWERA, Service de Chimie Quantique et Photophysique, C.P. 160/09, Université Libre de Bruxelles, 50 avenue F.D. Roosevelt, B-1050 Brussels, Belgium.

Recently, we have presented a tensorial formalism^{a}^{,}^{b} adapted to the spectroscopy of X_{2}Y_{4} molecules. It is based on formalisms already developed in Dijon for spherical-top molecules^{c} as well as on the work of Sartakov^{d}. This approach has the advantages to allow a systematic development of rovibrational interactions and to make global analyses easier to perform. We have used this tool to perform a re-analysis of the ν_{10}/ν_{7}/ν_{4}/ν_{12} tetrad both in frequencies and intensities in the 800–1500 cm^{−1} region. We have used 8943 infrared data in the ν_{10}/ν_{7}/ν_{4} region from W. E. Blass’ atlas^{e} completed by 5420 more recent data in the same range^{f} (207 of these having a precision of a few 10^{−6} cm^{−1}) and by 1138 data in the ν_{12} region from a spectrum recorded in Brussels. We obtain a global reduced RMS of 0.54 for line positions, which is comparable to the result of Ref. f. Concerning line intensities, we used 10657 data from Blass’ atlas and 25 high-precision diode laser data^{g}. We obtain a 11.6 % RMS in this case.

aW. Raballand, M. Rotger, V. Boudon and M. Loëte, J. Mol. Spectrosc. 217, 239–248 (2003).

bCh. Wenger, W. Raballand, M. Rotger and V. Boudon, J. Quant. Spectrosc. Radiat. Transfer 95, 521–538 (2005).

cV. Boudon, J.-P. Champion, T. Gabard, M. Loëte, F. Michelot, G. Pierre, M. Rotger, Ch. Wenger and M. Rey, J. Mol. Spectrosc.

228, 620–634 (2004).

dB. G. Sartakov, J. Oomens, J. Reuss and A. Fayt, J. Mol. Spectrosc. 185 31–47 (1997).

eW. E. Blass, J. J. Hillman, A. Fayt, S. J. Daunt, L. R. Senesac, A. C. Ewing, L. W. Jennings, J. S. Hager, S. L. Mahan, D. C.

Reuter and M. Sirota, J. Quant. Spectrosc. Radiat. Transfer 71, 47–60 (2001).

fF. Willaert, J. Demaison, L. Margulès, H. Mäder, H. Spahn, T. Giesen and A. Fayt, Mol. Phys. 104, 273–292 (2006).

gJ. Walrand, M. Lengelé, Gh. Blanquet and M. Lepère, Spectrochim. Acta A 59 421–426 (2003).